# Adaptive Trajectory Bundle Method for Roll-to-Roll Manufacturing Systems

### v1

@@ V-A Simulation Setup
where $E$ denotes Young’s modulus, $A$ the cross-sectional area, and $T_{i}(t)$ the tension setpoint for span $i$.
V-B Implementation Details
The Adaptive TBM was implemented with prediction horizon $H=15$ steps ($0.15$s). Following the formulation in Section III, at each timestep the controller constructs bundle matrices and solves the convex subproblem (12):
$\min_{\alpha,s,w,d^{\pm}}$ $\sum_{k=1}^{H}\|W_{r}^{(k)}\alpha^{(k)}\|_{2}^{2}+\mu\sum_{k=1}^{H-1}\left(\|s_{k}\|_{1}+\|w_{k}\|_{1}\right)$ (37)
$+\gamma_{+}\sum_{k=1}^{H-1}\mathbf{1}^{\top}d_{k}^{+}+\gamma_{-}\sum_{k=1}^{H-1}\mathbf{1}^{\top}d_{k}^{-}$
$W_{c,\mathrm{hard}}^{(k)}\alpha^{(k)}+w_{k}\geq 0,$
$W_{c,+}^{(k)}\alpha^{(k)}+d_{k}^{+}\geq 0,$
$W_{c,-}^{(k)}\alpha^{(k)}+d_{k}^{-}\geq 0,$
$\alpha^{(k)}\in\Delta^{m-1},\quad k=1,\ldots,H,$
where the cost residual bundles $W_{r}^{(k)}$ encode the quadratic tracking objectives:
$r_{k}(x_{k},u_{k})=\begin{bmatrix}Q^{1/2}(x_{k}-x_{k}^{r})\\ R^{1/2}(u_{k}-u_{k}^{r})\\ S^{1/2}\Delta u_{k}\end{bmatrix}$. (38)
The hard constraint bundles $W_{c,\mathrm{hard}}^{(k)}$ encode state and input bounds $(x_{\min}\leq x_{k}\leq x_{\max},u_{\min}\leq u_{k}\leq u_{\max})$, while soft constraint bundles $W_{c,+}^{(k)}$ and $W_{c,-}^{(k)}$ capture asymmetric tension bounds $(T_{\mathrm{ref}}-\delta^{-}\leq T(x_{k})\leq T_{\mathrm{ref}}+\delta^{+})$. The trajectory is recovered via (13): $x_{k}^{*}=W_{x}^{(k)}\alpha^{(k)*}$, $u_{k}^{*}=W_{u}^{(k)}\alpha^{(k)*}$.
Asymmetric penalties $\gamma_{+}=100$ and $\gamma_{-}=10$ prioritize over-tension avoidance (breakage risk) over under-tension (wrinkling), following the multiple penalty class extension in Remark 1. State weights $Q=\mathrm{diag}(100I_{N},10I_{N})$ emphasize tension regulation, while rate penalty $S=0.1I$ ensures smooth actuation.
The implementation follows Algorithm 1 with three key components:
Bundle Generation
At each time index $k$, $m=6N+21$ samples are generated via a hybrid strategy satisfying Assumption 3. A deterministic stencil of $6N$ samples perturbs the nominal $(\bar{x}_{k},\bar{u}_{k})$ by $\pm\Delta^{\ell}$ along each coordinate axis, ensuring sufficient directional coverage. An additional 20 stochastic samples drawn from $\mathcal{N}((\bar{x}_{k},\bar{u}_{k}),\Sigma)$ enhance convex hull coverage. The current iterate is always included in the sample set, as required by Assumption 3(ii).
Convex Subproblem
The bundle matrices $(W_{x}^{(k)},W_{u}^{(k)},W_{f}^{(k)},W_{r}^{(k)},W_{c}^{(k)})$ are constructed via (7)–(11), transforming the nonlinear R2R control problem into the convex subproblem (37) over simplex-constrained interpolation weights $\alpha^{(k)}$. As noted in Section III, the subproblem is convex: the objective combines a convex quadratic term with convex $\ell_{1}$ penalties, and all constraints are linear over the Cartesian product of simplices and non-negative orthants. The subproblem is solved using a standard convex optimization solver in 10–15 ms per iteration.
Adaptive Trust Region and Penalties
Following Definition 1, constraint violations are computed as $\nu_{\mathrm{dyn}}^{\ell}=\max_{k}\|s_{k}^{*}\|_{\infty}$ and $\nu_{\mathrm{hard}}^{\ell}=\max_{k}\|w_{k}^{*}\|_{\infty}$. The trust region adapts via (17) with expansion factor $\beta_{\exp}=1.5$, contraction factor $\beta_{\mathrm{con}}=0.5$, feasibility threshold $\tau_{\mathrm{feas}}=10^{-4}$, and violation threshold $\tau_{\mathrm{viol}}=10^{-2}$. Penalties adapt via (18)–(19) with increase factor $\rho_{\mu}=\rho_{\gamma}=2.0$. This enables robust handling of aggressive transients without manual tuning, as guaranteed by Theorem 1.
V-C Tension Step Tracking
@@
V-D Velocity Setpoint Change
@@
V-E Convergence Comparison
Figure 4 compares the convergence behavior of TBM and Adaptive TBM. The adaptive variant reduces peak transient error by 7.4% (3.15 N vs 3.4 N) in tension step tracking, as the trust region mechanism constrains step sizes during large disturbances. Additionally, Adaptive TBM converges approximately 50 iterations faster, benefiting from trust region expansion during refinement. Both methods achieve comparable steady-state accuracy (1.4 N and 0.8 N for the two scenarios), demonstrating that the adaptive mechanisms improve transient response without sacrificing final performance.
Figure 4: Convergence comparison between TBM and Adaptive TBM.
@@ VI Conclusion
This paper developed an adaptive trajectory bundle method for R2R manufacturing control. The proposed framework combines the derivative-free simplicity of sampling-based methods with the constraint-handling rigor of sequential convex programming. Adaptive trust region and penalty mechanisms eliminate manual tuning while ensuring robust convergence. Simulation results on a six-zone R2R system demonstrate tracking accuracy comparable to gradient-based MPC and superior disturbance rejection over MPPI baselines. The approach offers a practical solution for industrial web handling where analytical gradients are unavailable or computationally prohibitive.

### v2

@@ IV-A Implementation Details
The Adaptive TBM was implemented with prediction horizon $H=15$ steps ($0.15$s), solving the convex subproblem (12) at each timestep. The cost residual encodes quadratic tracking:
$r_{k}(x_{k},u_{k})=\begin{bmatrix}Q^{1/2}(x_{k}-x_{k}^{r})\\ R^{1/2}(u_{k}-u_{k}^{r})\\ S^{1/2}\Delta u_{k}\end{bmatrix}$, (21)
with $Q=\mathrm{diag}(100I_{N},10I_{N})$, $R=I$, and $S=0.1I$. Hard constraints enforce state/input bounds, while soft constraints capture asymmetric tension limits with penalties $\gamma_{+}=100$ (over-tension) and $\gamma_{-}=10$ (under-tension) to prioritize breakage avoidance over wrinkling.
At each time index $k$, $m=6N+21$ samples are generated: a deterministic stencil of $6N$ samples perturbing $(\bar{x}_{k},\bar{u}_{k})$ by $\pm\Delta^{\ell}$ along coordinate axes, plus 20 stochastic samples from $\mathcal{N}((\bar{x}_{k},\bar{u}_{k}),\Sigma)$. The resulting convex subproblem is solved in 10–15 ms. Trust region parameters are $\beta_{\exp}=1.5$, $\beta_{\mathrm{con}}=0.5$, $\tau_{\mathrm{feas}}=10^{-4}$, and $\tau_{\mathrm{viol}}=10^{-2}$; penalties adapt with $\rho_{\mu}=\rho_{\gamma}=2.0$.
IV-B Tension Step Tracking
@@
IV-C Velocity Setpoint Change
@@
V Experimental Validation: R2R Dry Transfer
This section validates the proposed Adaptive TBM on an R2R mechanical dry transfer system for transferring advanced 2D materials.
V-A Dry Transfer Dynamics
The R2R dry transfer process, illustrated in Fig. 5, peels functional 2D material from a donor substrate and deposits it onto a target substrate. The system comprises three driven rollers: one unwinding (roller 1) and two rewinding (rollers 2 and 3).
Figure 5: Illustration of the R2R peeling front.
When peeling occurs, the energy balance at the peeling front yields [zhao2022dynamic, hong2023adhesion]:
$T_{2}+T_{3}-T_{1}=\sigma,$ (22)
where $T_{i}$ denotes web tension in section $i$, and $\sigma$ combines bending and adhesion energy effects. The web velocity dynamics follow from torque balance on the rewinding rollers:
$\dot{v}_{i}=-\frac{R_{i}^{2}}{J_{i}}T_{i}+\frac{R_{i}}{J_{i}}u_{i}-\frac{f_{i}}{R_{i}}v_{i},\quad i=2,3,$ (23)
where $R_{i}$, $J_{i}$, $f_{i}$ are the radius, inertia, and friction coefficient of roller $i$, and $u_{i}$ is the motor torque input. The unstretched web length dynamics follow from mass conservation:
$\dot{l}_{1}=\frac{v_{1}-v_{p}}{1+\varepsilon_{1}},\quad\dot{l}_{i}=\frac{v_{p}}{1+\varepsilon_{1}}-\frac{v_{i}}{1+\varepsilon_{i}},\quad i=2,3,$ (24)
where $\varepsilon_{i}=T_{i}/(A_{i}E_{i})$ is strain and $v_{p}$ is the peeling front velocity. Using the one-to-one relationship between tensions and unstretched lengths [zhao2022dynamic], the tension derivatives are:
$\dot{T}_{i}=\frac{\partial T_{i}}{\partial l_{1}}\dot{l}_{1}+\frac{\partial T_{i}}{\partial l_{2}}\dot{l}_{2}+\frac{\partial T_{i}}{\partial l_{3}}\dot{l}_{3},\quad i=2,3.$ (25)
Equations (22)–(25) define the state-space representation:
$\dot{x}=f(x,w,u),\quad x=[v_{2},v_{3},T_{2},T_{3}]^{\top},\quad u=[u_{2},u_{3}]^{\top},$ (26)
where $w=[v_{p},\sigma]^{\top}$ contains the exogenous inputs. Since $\sigma$ varies periodically with the material pattern and $v_{p}$ is a nonlinear function of system states, both are treated as bounded disturbances.
V-B Experimental Setup
[…]
Figure 6: The experimental testbed.
The control hardware employs a National Instruments CompactRIO to collect data from quadrature encoders on rollers 2 and 3, and load cells on each web section.
[…]
Three controllers are compared: (i) LQR [martin2025nldi], (ii) TBM with fixed penalties [tracy2025trajectory], and (iii) the proposed Adaptive TBM. The LQR controller was tuned using Bryson’s rule. All methods use identical cost weights and constraints.
V-C Results
Two test cases evaluate the controllers across the operating envelope: Case 1 with tensions stepping from 10 N to 18 N at $v_{1}=0.005$ m/s, and Case 2 with tensions stepping from 14 N to 22 N at $v_{1}=0.011$ m/s. The step command is applied at $t=5$ s.
Figure 7 shows the tension responses for both cases. The LQR controller (blue) tracks the reference but exhibits a slower response, requiring approximately 4–5 s to fully settle. Both TBM variants achieve faster transient response than LQR. The standard TBM (red) shows moderate overshoot, particularly visible in $T_{2}$ for Case 2. The Adaptive TBM (purple) demonstrates improved performance, achieving faster settling with reduced overshoot through automatic trust region adjustment during transients. Across both cases, the Adaptive TBM consistently outperforms the standard TBM and LQR.
Figure 7: Experimental tension responses for Case 1 and Case 2 . Top row: $T_{2}$; bottom row: $T_{3}$.
@@ VI Conclusion
This paper developed an adaptive trajectory bundle method for constrained R2R control, combining derivative-free optimization with rigorous constraint handling. The method approximates nonlinear dynamics through bundle interpolation and employs adaptive trust region and penalty mechanisms that eliminate manual tuning. Simulations on a six-zone R2R system showed 4.3% and 11.1% lower tension RMSE than gradient-based MPC in step tracking and velocity transients, respectively, with smoother actuation than MPPI variants. Experimental validation on an R2R dry transfer system confirmed faster settling and reduced overshoot compared to LQR and non-adaptive TBM. The approach offers a practical solution for industrial web handling where gradients are unavailable and hard constraint satisfaction is critical.